# DCNv3: Towards Next Generation Deep Cross Network for Click-Through Rate Prediction

### v4

@@ Abstract.
Deep & Cross Network and its derivative models have become an important paradigm in click-through rate (CTR) prediction due to their effective balance between computational cost and performance. However, these models face four major limitations: (1) while most models claim to capture high-order feature interactions, they often do so implicitly and non-interpretably through deep neural networks (DNN), which limits the trustworthiness of the model’s predictions; (2) the performance of existing explicit feature interaction methods is often weaker than that of implicit DNN, undermining their necessity; (3) many models fail to adaptively filter noise while enhancing the order of feature interactions; (4) the fusion methods of most models cannot provide suitable supervision signals for their different interaction methods.
To address the identified limitations, this paper proposes the next generation Deep Cross Network (DCNv3) and Shallow & Deep Cross Network (SDCNv3). These models ensure interpretability in feature interaction modeling while exponentially increasing the order of feature interactions to achieve genuine Deep Crossing rather than just Deep & Cross. Additionally, we employ a Self-Mask operation to filter noise and reduce the number of parameters in the cross network by half. In the fusion layer, we use a simple yet effective loss weight calculation method called Tri-BCE to provide appropriate supervision signals. Comprehensive experiments on six datasets demonstrate the effectiveness, efficiency, and interpretability of DCNv3 and SDCNv3. The code, running logs, and detailed hyperparameter configurations are available at: https://anonymous.4open.science/r/DCNv3-E352.
@@ 1. Introduction
  • Lack of interpretability. As shown in Figure 1, most models integrate deep neural networks (Covington et al., 2016) (DNN) to model implicit high-order feature interactions and achieve AUC performance between 81.3 and 81.5. This demonstrates the effectiveness of implicit feature interactions. However, implicit feature interactions lack interpretability (Marcus, 2018), which significantly reduces the trustworthiness of deep CTR model predictions.
  • Limited necessity for explicit interactions. As observed in Figure 1, most models (e.g., CIN (Lian et al., 2018)) using only explicit feature interactions achieve AUC performance below 81.3, while DNN achieves a better AUC of 81.4. This indicates that the performance of most explicit modeling methods is weaker than that of implicit DNN, which undoubtedly undermines the necessity of integrating explicit and implicit feature interactions. Therefore, FinalMLP (Mao et al., 2023) attempts to model features implicitly in a dual manner, discarding traditional explicit interaction methods, and thereby achieving state-of-the-art performance.
  • Ineffective noise filtering capability. Many studies (Liu et al., 2020; Zhu et al., 2023a; Deng et al., 2021; Wang et al., 2023a) point out that CTR models contain a significant amount of redundant feature interactions and noise, especially in higher-order feature interactions. Consequently, most CTR models (Song et al., 2019; Wang et al., 2021b; Lian et al., 2018) are built with only two or three network layers, abandoning the explicit capture of effective higher-order feature interaction information. Meanwhile, filtering noise for the model often incurs additional computational costs, which can lead to longer training and inference times, potentially offsetting the benefits gained from improved model accuracy.
@@
To address the aforementioned limitations, this paper proposes the next generation Deep Cross Network (DCNv3) and the Shallow & Deep Cross Network (SDCNv3), which integrates both low-order and high-order feature interactions while ensuring model interpretability by avoiding the use of DNN for implicit high-order feature interaction modeling. Specifically, we introduce a new exponentially growing Deep Crossing method to explicitly model high-order feature interactions, and we design a Self-Mask operation to filter noise and reduce the number of parameters in the cross network by half. In the fusion layer, we propose a simple yet effective multi-loss trade-off strategy and calculation method, called Tri-BCE, to provide suitable supervision signals for different sub-networks.
@@
  • To the best of our knowledge, this is the first work to achieve surprising performance using only explicit feature interaction modeling without integrating DNN, which may contrast with the popular paradigms in the past CTR prediction literature.
@@ 2.1. CTR Prediction
Effectively capturing feature interactions has always been one of the key methods for improving CTR prediction, thus receiving extensive research attention (Tian et al., 2023; Wang et al., 2023a; Zhu et al., 2023b). Traditional methods include LR (Richardson et al., 2007), which captures first-order feature interactions, and FM (Rendle, 2010) and its derivatives (Sun et al., 2021; Xiao et al., 2017; Pan et al., 2018), which capture second-order feature interactions. With the rise of deep learning, several models attempt to use DNN to capture higher-order feature interactions (e.g., PNN (Qu et al., 2016), Wide & Deep (Cheng et al., 2016), DeepFM (Guo et al., 2017), DCNv1 (Wang et al., 2017), DCNv2 (Wang et al., 2021b), and DIN (Zhou et al., 2018)), achieving better performance. Among these, the DCN series models are widely recognized for their effective trade-off between efficiency and performance, gaining significant attention from both academia and industry (Wang et al., 2017, 2021b; Chen et al., 2021; Wang et al., 2023a; Yu et al., 2021; Lian et al., 2018). Most subsequent deep CTR models follow the paradigm established by DCN, integrating explicit and implicit feature interactions.
Explicit feature interactions are often modeled directly through hierarchical structures, such as the Cross Layer in the DCN (Wang et al., 2017), the Graph Layer in FiGNN (Li et al., 2019), and the Interacting Layer in AutoInt (Song et al., 2019). These methods ensure partial interpretability while allowing the capture of higher-order feature interactions. On the other hand, implicit feature interactions use DNN to automatically learn complex, non-manually defined patterns and interactions among features, further enhancing model performance (Wang et al., 2021a; Cheng et al., 2020; Wang et al., 2021b).
@@ 2.2. Problem Definition
Explicit feature interaction aims to model the combinations and relationships between input features directly through predefined functions, thereby improving model interpretability. A popular explicit feature interaction method is $X_{2}=X_{1}\odot X_{1}$ (Wang et al., 2021a, 2017; He and Chua, 2017; Wang et al., 2021b; Chen et al., 2021). This method uses the Hadamard Product to interact with two first-order features and generate a second-order feature.
@@ 3.1. Embedding & Reshape Layer
The input $X$ of the CTR prediction task, which is multi-field categorical data and is represented using one-hot encoding. Most CTR prediction models (Song et al., 2019; Li et al., 2023; Sang et al., 2024) utilize an embedding layer to transform them into low-dimensional dense vectors: $\mathbf{e}_{i}=\textit{E}_{i}x_{i}$, where $\textit{E}_{i}\in\mathbb{R}^{d\times s_{i}}$ and $s_{i}$ separately indicate the embedding matrix and the vocabulary size for the $i$-th field, $d$ represents the embedding dimension. In our model, to enable the Cross & Masked Vector to share a weight matrix and separately process two different information streams, we reshape the embeddings using the chunk operation to divide them into two different views: $\mathbf{e}_{i,a},\mathbf{e}_{i,b}=\texttt{chunk}(\mathbf{e}_{i})$. Further, we can get the resulting embedding after reshaping: $\mathbf{x_{1}}=\left[\mathbf{e}_{1,a},\cdots,\mathbf{e}_{f,a},\mathbf{e}_{1,b},\cdots,\mathbf{e}_{f,b}\right]\in\mathbb{R}^{D}$, where $f$ denotes the number of fields, $D=\sum_{i=1}^{f}d$, and $\mathbf{x_{1}}$ denotes the first-order features.
@@ 3.2. Shallow & Deep Cross Network v3
The architecture of SDCNv3, shown in Figure 2, integrates different Crossing methods and various layers to simultaneously capture both high-order and low-order explicit feature interactions. As the core structure of SDCNv3, the shallow explicit modeling of SCNv3 recursive formula is as follows:
@@
where $\mathbf{c}_{l}\in\mathbb{R}^{\frac{D}{2}}$ represents the Cross Vector at $l\text{-}th$ layer, Mask denotes the Self-Mask operation, $\mathbf{W}_{l}\in\mathbb{R}^{\frac{D}{2}\times D}$ and $\mathbf{b}_{l}\in\mathbb{R}^{\frac{D}{2}}$ are the learnable weight matrix and bias vector, respectively, and $\mathbf{x}_{l}\in\mathbb{R}^{D}$ represents the $l\text{-}th$ order feature interaction.
@@
The deep explicit modeling of DCNv3 recursive formula is as follows:
@@
Figure 3 visualizes the computation process of different crossing methods. The Cross Vector $\mathbf{c}_{l}$ aims to model interactions between features at the bit-level, while the weight matrix $\mathbf{W}_{l}$ aims to compute the inherent importance of different feature fields. However, as pointed out by some works (Xiao et al., 2017; Liu et al., 2020; Wang et al., 2023a), not all feature interactions are beneficial for the final prediction in CTR tasks. Therefore, we introduce the Self-Mask operation to filter out noisy information from feature interactions in another view while maintaining the integrity of the interaction information of the original view. To avoid additional computational costs, we use LayerNorm (Ba et al., 2016) with element-wise affine to regularize it, ensuring a mask rate of around 0.5. The specific formalization of the Self-Mask operation is as follows:
(4) $\texttt{Mask}(\mathbf{c}_{l})=\mathbf{c}_{l}\odot\texttt{max}(0,\texttt{LayerNorm}(\mathbf{c}_{l})).$
Other mask mechanisms can also be used here, such as random Mask based on the Bernoulli distribution, learnable Mask based on Top-K selection, etc. To ensure our proposed model is simple and effective, we use LayerNorm (Ba et al., 2016) to perform a straightforward and efficient normal distribution transformation on $\mathbf{c}_{l}$, ensuring that its output contains approximately 50% zero values to filter out noise, and it reduces the number of parameters by half compared to Cross Network v2.
3.3. Fusion Layer
Most previous CTR models attempt to capture both implicit and explicit feature interactions, which essentially means capturing low-order and high-order feature interactions. Our SDCNv3 achieves this only through explicit modeling. On the other hand, existing works (Wang et al., 2021b, 2023a; Zhu et al., 2023b) often integrate sub-networks using either parallel or stacked structures. Considering the high parallelism of the former, we use a parallel structure to fuse information and compute the loss:
@@ 3.3. Fusion Layer
where $\mathbf{W}_{\scriptscriptstyle D}$ and $\mathbf{W}_{\scriptscriptstyle S}\in\mathbb{R}^{1\times D}$ represent learnable weights, $\mathbf{b}_{\scriptscriptstyle D}$ and $\mathbf{b}_{\scriptscriptstyle S}$ are biases, Mean denotes the mean operation, $\hat{y}_{\scriptscriptstyle D},\hat{y}_{\scriptscriptstyle S}$ represent the prediction results of DCNv3 and SCNv3, respectively, and $L$ denotes the last number of layers.
@@
As demonstrated by (Li et al., 2024), providing a single supervision signal to sub-networks is often suboptimal. Our proposed Tri-BCE loss helps sub-networks learn better parameters by providing adaptive weights that change throughout the learning process. Theoretically, we can derive the gradients obtained by $\hat{y}_{\scriptscriptstyle D}$:
@@ 3.4. Complexity Analysis
To further compare the time complexity of the DCN series models, we discuss and analyze the time complexity of different models. Let $W_{\Psi}$ denote the predefined number of parameters in the DNN, and $s$ denote the feature vocabulary size. The definitions of the other variables can be found in the previous sections. For clarity, we further provide a comparison of the magnitudes of different variables in Table 1. We can derive:
@@ 4.1. Experiment Setup
We compared DCNv3 and SDCNv3 with some state-of-the-art (SOTA) models (* denotes Integrating the original model with DNN networks): (1) Since DCNv3 is a standalone network that performs explicit feature interactions, we compare it with several models that also perform explicit feature interactions on two large-scale datasets. For example, LR (2007) (Richardson et al., 2007) implements first-order feature interactions; FM and its derivative models FM (2010) (Rendle, 2010), FwFM (2018) (Pan et al., 2018), AFM (2017) (Xiao et al., 2017), FmFM (2021) (Sun et al., 2021) implement second-order feature interactions; and CrossNetv1 (2017) (Wang et al., 2017), CrossNetv2 (2021) (Wang et al., 2021b), CIN (2018) (Lian et al., 2018), AutoInt (2019) (Song et al., 2019), AFN (2020) (Cheng et al., 2020), FiGNN (2019) (Li et al., 2019) implement higher-order feature interactions. (2) To verify the superiority of DCNv3 and SDCNv3 over models that include implicit feature interactions, we further select several high-performance representative baselines, such as PNN (2016) (Qu et al., 2016), Wide & Deep (2016) (Cheng et al., 2016), DeepFM (2017) (Guo et al., 2017), DCNv1 (2017) (Wang et al., 2017), xDeepFM (2018) (Lian et al., 2018), AutoInt* (2019) (Song et al., 2019), AFN* (2020) (Cheng et al., 2020), DCNv2 (2021) (Wang et al., 2021b), EDCN (2021) (Chen et al., 2021), MaskNet (2021) (Wang et al., 2021a), CL4CTR (2023) (Wang et al., 2023b), EulerNet (2023) (Tian et al., 2023), FinalMLP (2023) (Mao et al., 2023), and FINAL (2023) (Zhu et al., 2023b).

### v5

@@ Abstract.
Deep & Cross Network and its derivative models have become an important paradigm for click-through rate (CTR) prediction due to their effective balance between computational cost and performance. However, these models face four major limitations: (1) the performance of existing explicit feature interaction methods is often weaker than that of implicit DNN, undermining their necessity; (2) many models fail to adaptively filter noise while enhancing the order of feature interactions; (3) the fusion methods of most models cannot provide suitable supervision signals for their different interaction procedures; (4) while most models claim to capture high-order feature interactions, they often do so implicitly and non-interpretably through deep neural networks (DNN), which limits the trustworthiness of the model’s predictions.
To address the identified limitations, this paper proposes the next generation Deep Cross Network (DCNv3) and Shallow & Deep Cross Network (SDCNv3) for CTR prediction. These models ensure interpretability in feature interaction modeling while exponentially increasing the order of feature interactions to achieve genuine Deep Crossing rather than just Deep & Cross. Additionally, we employ a Self-Mask operation to filter noise and reduce the number of parameters in the cross network by half. In the fusion layer, we use a simple yet effective loss weight calculation method called Tri-BCE to provide appropriate supervision signals. Comprehensive experiments on six datasets demonstrate the effectiveness, efficiency, and interpretability of DCNv3 and SDCNv3. The code, running logs, and detailed hyperparameter configurations are available at: https://github.com/salmon1802/DCNv3.
@@ 1. Introduction
  • Limited necessity for explicit interactions. As observed in Figure 1, most models (e.g., CIN (Lian et al., 2018)) using only explicit feature interactions achieve AUC performance below 81.3, while deep neural network (Covington et al., 2016) (DNN) achieves a better AUC of 81.4. This indicates that the performance of most explicit modeling methods is weaker than that of implicit DNN, which undoubtedly undermines the necessity of integrating explicit feature interactions. However, some works (Rendle et al., 2020; Wang et al., 2017) also highlight the limitations of DNN in capturing multiplicative feature interactions. Therefore, exploring a more effective method for explicit feature interactions is necessary. FINAL (Zhu et al., 2023b) is the first model to employ an exponentially growing feature interaction method, incorporating the idea of residual connections (He et al., 2016) to capture both explicit and implicit feature interactions within the linear layers of DNN, thereby achieving state-of-the-art (SOTA) performance.
  • Ineffective noise filtering capability. Many studies (Liu et al., 2020; Zhu et al., 2023a; Deng et al., 2021; Wang et al., 2023a) point out that CTR models contain a significant amount of redundant feature interactions and noise, especially in high-order feature interactions. Consequently, most CTR models (Song et al., 2019; Wang et al., 2021b; Lian et al., 2018) are built with only two or three network layers, abandoning the explicit capture of effective high-order feature interaction information. Meanwhile, filtering noise for the model often incurs additional computational costs, which can lead to longer training and inference times, potentially offsetting the benefits gained from improved model accuracy.
@@
  • Lack of interpretability. As shown in Figure 1, most models integrate DNN to model implicit high-order feature interactions and achieve AUC performance between 81.3 and 81.5. This demonstrates the effectiveness of implicit feature interactions. However, implicit feature interactions lack interpretability (Marcus, 2018), which reduces the trustworthiness of CTR model predictions.
To address the aforementioned limitations, this paper proposes the next generation Deep Cross Network (DCNv3) and the Shallow & Deep Cross Network (SDCNv3), which integrates both low-order and high-order feature interactions while ensuring model interpretability by avoiding the use of DNN for implicit high-order feature interaction modeling. Specifically, we introduce a new exponentially growing Deep Crossing method to explicitly model high-order feature interactions, and we design a Self-Mask operation to filter noise and reduce the number of parameters in the cross network by half. In the fusion layer, we propose a simple yet effective multi-loss calculation method, called Tri-BCE, to provide suitable supervision signals for different sub-networks.
@@
  • To the best of our knowledge, this is the first work to achieve surprising performance using only explicit feature interaction modeling without integrating DNN, which may contrast with the popular paradigms in the past CTR prediction literature. Therefore, this work may inspire further revisiting and innovation in feature interaction modeling.
@@ 2.1. CTR Prediction
Effectively capturing feature interactions has always been one of the key methods for improving CTR prediction, thus receiving extensive research attention (Tian et al., 2023; Wang et al., 2023a; Zhu et al., 2023b). Traditional methods include LR (Richardson et al., 2007), which captures first-order feature interactions, and FM (Rendle, 2010) and its derivatives (Sun et al., 2021; Xiao et al., 2017; Pan et al., 2018), which capture second-order feature interactions. With the rise of deep learning, several models attempt to use DNN to capture high-order feature interactions (e.g., PNN (Qu et al., 2016), Wide & Deep (Cheng et al., 2016), DeepFM (Guo et al., 2017), DCNv1 (Wang et al., 2017), DCNv2 (Wang et al., 2021b), and DIN (Zhou et al., 2018)), achieving better performance. Among these, the DCN series models are widely recognized for their effective trade-off between efficiency and performance, gaining significant attention from both academia and industry (Wang et al., 2017, 2021b; Chen et al., 2021; Wang et al., 2023a; Yu et al., 2021; Lian et al., 2018). Most subsequent deep CTR models follow the paradigm established by DCN, integrating explicit and implicit feature interactions.
Explicit feature interactions are often modeled directly through hierarchical structures, such as the Cross Layer in the DCN (Wang et al., 2017), the Graph Layer in FiGNN (Li et al., 2019), and the Interacting Layer in AutoInt (Song et al., 2019). These methods ensure partial interpretability while allowing the capture of high-order feature interactions. On the other hand, implicit feature interactions use DNN to automatically learn complex, non-manually defined patterns and interactions among features, further enhancing model performance (Wang et al., 2021a; Cheng et al., 2020; Wang et al., 2021b).
@@ 2.2. Problem Definition
Explicit feature interaction aims to model the combinations and relationships between input features directly through predefined functions, thereby improving model interpretability. A popular method for linearly increasing explicit feature interaction is $X_{n}=X_{1}\odot X_{n-1}$ (Wang et al., 2017, 2021b, 2023a). This method uses the Hadamard Product to interact feature $X_{n-1}$ with $X_{1}$ to generate the n-$th$ order feature $X_{n}$. Meanwhile, an exponentially growing feature interaction method is $X_{2^{n}}=X_{2^{n-1}}\odot X_{2^{n-1}}$, which can generate high-order features more efficiently.
@@ 3. Proposed Architecture
The architecture of SDCNv3, shown in Figure 2, integrates different explicit feature interaction methods: SCNv3 and DCNv3, enabling it to simultaneously capture both low-order and high-order feature interactions while ensuring the model’s interpretability.
@@ 3.1. Embedding & Reshape Layer
The input $X$ of the CTR prediction task, which is multi-field categorical data and is represented using one-hot encoding. Most CTR prediction models (Song et al., 2019; Li et al., 2023; Sang et al., 2024) utilize an embedding layer to transform them into low-dimensional dense vectors: $\mathbf{e}_{i}=\textit{E}_{i}x_{i}$, where $\textit{E}_{i}\in\mathbb{R}^{d\times s_{i}}$ and $s_{i}$ separately indicate the embedding matrix and the vocabulary size for the $i$-th field, $d$ represents the embedding dimension. In our model, to enable the network to obtain richer information from multi-views, we reshape the embeddings using the chunk operation (Li et al., 2024a) to divide them into two different views: $\mathbf{e}_{i,a},\mathbf{e}_{i,b}=\texttt{chunk}(\mathbf{e}_{i})$, where $\mathbf{e}_{i,a}$ is considered an original view, and $\mathbf{e}_{i,b}$ is considered another view. Further, we can get the resulting embedding after reshaping: $\mathbf{x_{1}}=\left[\mathbf{e}_{1,a},\cdots,\mathbf{e}_{f,a},\mathbf{e}_{1,b},\cdots,\mathbf{e}_{f,b}\right]\in\mathbb{R}^{D}$, where $f$ denotes the number of fields, $D=\sum_{i=1}^{f}d$, and $\mathbf{x_{1}}$ denotes the first-order features shared as input by SCNv3 and DCNv3.
@@ 3.2. Shallow & Deep Cross Network v3
SCNv3: It is used for low-order (shallow) explicit feature interaction with linear growth. Its recursive formula is as follows:
@@
where $\mathbf{c}_{l}\in\mathbb{R}^{\frac{D}{2}}$ represents the Cross Vector at $l\text{-}th$ layer, Mask denotes the Self-Mask operation, $\mathbf{W}_{l}\in\mathbb{R}^{\frac{D}{2}\times D}$ and $\mathbf{b}_{l}\in\mathbb{R}^{\frac{D}{2}}$ are the learnable weight matrix and bias vector, respectively, $\mathbf{x}_{l}\in\mathbb{R}^{D}$ represents the $l\text{-}th$ order feature interaction, and Mask represents the Self-Mask, which will be detailed later.
@@
DCNv3: As the core structure of SDCNv3, it is used for high-order (deep) explicit feature interaction with exponential growth. Its recursive formula is as follows:
@@
Self-Mask: Figure 3 visualizes the computation process of different crossing methods. The Cross Vector $\mathbf{c}_{l}$ aims to model interactions between features at the bit-level, while the weight matrix $\mathbf{W}_{l}$ aims to compute the inherent importance of different feature fields. However, as pointed out by some works (Xiao et al., 2017; Liu et al., 2020; Wang et al., 2023a), not all feature interactions are beneficial for the final prediction in CTR tasks. Therefore, we introduce the Self-Mask operation to filter out noisy information from feature interactions in another view (corresponding to $\texttt{Mask}(\mathbf{c}_{l})$) while maintaining the integrity of the interaction information of the original view (corresponding to $\mathbf{c}_{l}$). The Self-Mask operation is as follows:
(4) $\texttt{Mask}(\mathbf{c}_{l})=\mathbf{c}_{l}\ \odot\ \texttt{max}\left(0,\ \texttt{LN}(\mathbf{c}_{l})\right),$
$\texttt{LN}(\mathbf{c}_{l})=\mathbf{g}\odot\text{Norm}(\mathbf{c}_{l})+\mathbf{b},\quad\text{Norm}(\mathbf{c}_{l})=\frac{\mathbf{c}_{l}-\mu}{\delta}$
$\mu=\frac{2}{D}\sum_{i=1}^{\frac{D}{2}}\mathbf{c}_{l,i},\quad\delta=\sqrt{\frac{2}{D}\sum_{i=1}^{\frac{D}{2}}\left(\mathbf{c}_{l,i}-\mu\right)^{2}},$
where LN denotes LayerNorm (Ba et al., 2016), $\mu$ and $\delta$ are the mean and standard deviation of $\mathbf{c}_{l}$, the gain $\mathbf{g}$ and bias $\mathbf{b}$ are parameters of the same dimension, and $\mathbf{c}_{l,i}$ denotes the i-$th$ element in $\mathbf{c}_{l}$. Other mask mechanisms can also be used here, such as random Mask based on the Bernoulli distribution, learnable Mask based on Top-K selection, etc. To ensure our proposed model is simple yet effective, we use LayerNorm to perform a normal distribution transformation on $\mathbf{c}_{l}$, ensuring that $\texttt{Mask}(\mathbf{c}_{l})$ contains approximately 50% zero values to filter out noise and improve computational efficiency.
Fusion: Most previous CTR models attempt to capture both implicit and explicit feature interactions, which essentially means capturing low-order and high-order feature interactions. Our SDCNv3 achieves this only through explicit modeling. On the other hand, existing works (Wang et al., 2021b, 2023a; Zhu et al., 2023b) often integrate sub-networks using either parallel or stacked structures. Considering the high parallelism of the former, we use a parallel structure to fuse information:
@@
where $\mathbf{W}_{\scriptscriptstyle D}$ and $\mathbf{W}_{\scriptscriptstyle S}\in\mathbb{R}^{1\times D}$ represent learnable weights, $\mathbf{b}_{\scriptscriptstyle D}$ and $\mathbf{b}_{\scriptscriptstyle S}$ are biases, Mean denotes the mean operation, $\hat{y}_{\scriptscriptstyle D},\hat{y}_{\scriptscriptstyle S}$, $\hat{y}$ represent the prediction results of DCNv3, SCNv3, and SDCNv3, respectively, and $L$ denotes the last number of layers.
@@ 3.3. Tri-BCE Loss
as demonstrated by (Li et al., 2024b), providing a single supervision signal to sub-networks is often suboptimal. Our proposed Tri-BCE loss helps sub-networks learn better parameters by providing adaptive weights that change throughout the learning process. Theoretically, we can derive the gradients obtained by $\hat{y}_{\scriptscriptstyle D}$:
@@ 3.4. Complexity Analysis
To further compare the time complexity of the DCN series models, we discuss and analyze the time complexity of different models. Let $W_{\Psi}$ denote the predefined number of parameters in the DNN. The definitions of the other variables can be found in the previous sections. For clarity, we further provide a comparison of the magnitudes of different variables in Table 1. We can derive:
@@ 4.1. Experiment Setup
We compared DCNv3 and SDCNv3 with some SOTA models (* denotes Integrating the original model with DNN networks): (1) Since DCNv3 is a standalone network that performs explicit feature interactions, we compare it with several models that also perform explicit feature interactions on two large-scale datasets. For example, LR (2007) (Richardson et al., 2007) implements first-order feature interactions; FM and its derivative models FM (2010) (Rendle, 2010), FwFM (2018) (Pan et al., 2018), AFM (2017) (Xiao et al., 2017), FmFM (2021) (Sun et al., 2021) implement second-order feature interactions; and CrossNetv1 (2017) (Wang et al., 2017), CrossNetv2 (2021) (Wang et al., 2021b), CIN (2018) (Lian et al., 2018), AutoInt (2019) (Song et al., 2019), AFN (2020) (Cheng et al., 2020), FiGNN (2019) (Li et al., 2019) implement high-order feature interactions. (2) To verify the superiority of DCNv3 and SDCNv3 over models that include implicit feature interactions, we further select several high-performance representative baselines, such as PNN (Qu et al., 2016) and Wide & Deep (Cheng et al., 2016) (2016); DeepFM (Guo et al., 2017) and DCNv1 (Wang et al., 2017) (2017); xDeepFM (2018) (Lian et al., 2018); AutoInt* (2019) (Song et al., 2019); AFN* (2020) (Cheng et al., 2020); DCNv2 (Wang et al., 2021b) and EDCN (Chen et al., 2021), MaskNet (Wang et al., 2021a) (2021); CL4CTR (Wang et al., 2023b), EulerNet (Tian et al., 2023), FinalMLP (Mao et al., 2023), and FINAL (Zhu et al., 2023b) (2023).